# Cost Efficient Fairness Audit Under Partial Feedback

### v1

@@ 3 Our Results
Let us first define certain quantities that feature in our guarantees. We denote by $\beta\coloneqq\mathbb{P}[f=0],\beta_{a}\coloneqq\mathbb{P}[f=0\mid A=a],p_{y,a}\coloneqq\mathbb{P}[f=1,Y=y,A=a]$, $q_{y,a}\coloneqq\mathbb{P}[Y=y,A=a]$, $p_{y\mid a}\coloneqq\mathbb{P}[f=1,Y=y\mid A=a]$, $q_{y\mid a}\coloneqq\mathbb{P}[Y=y\mid A=a]$ for all $y\in\{0,1\}$ and $a\in{\mathcal{A}}$, and $\gamma_{ij}\coloneqq\mathbb{P}[f=i,Y=j]$ for $i,j\in\{0,1\}$. Note that $p_{y,a}=p_{y\mid a}\cdot\mathbb{P}[A=a]$, therefore $p_{y,a}\leq p_{y\mid a}$, and similarly, $q_{y,a}\leq q_{y\mid a}$. Next, we state the key results of this work.
@@
Given $\varepsilon,\delta\in(0,1)$, under the blackbox model, the audit policy that requests true labels of every individual with $f=0$ (Algorithm 3 in Appendix A) succeeds in the $(\varepsilon,\delta)$-fairness audit with $\texttt{cost}\leq\widetilde{O}\left(\sum_{a\in\mathcal{A}}\sum_{y\in\{0,1\}}\frac{c_{feat}\beta+c_{lab}\gamma_{00}}{q_{y,a}\varepsilon^{2}}\right)~.$ Moreover, the number of true labels requested by this policy is at most $\widetilde{O}(\sum_{a\in{\mathcal{A}}}\sum_{y\in\{0,1\}}\frac{\beta}{q_{y,a}\varepsilon^{2}})$.
Next, we give an algorithm RS-Audit (Algorithm 1) that has improved audit cost using rejection sampling. This algorithm requests an optimal number of true labels, indicating that characterization of the blackbox case is essentially complete. In the design of RS-Audit (as well as the natural baseline), we use a threshold-based stopping time construction to decide when to stop requesting true labels. The analysis requires concentration properties of the negative binomial distribution, instead of the usual Chernoff bounds (Lemma 28). Below we formally state the guarantees (proof in A.2).
@@
For $\varepsilon,\delta\in(0,1)$, in the blackbox model, Algorithm 1 (RS-Audit) succeeds in $(\varepsilon,\delta)$-fairness audit with $\texttt{cost}\leq\widetilde{O}\left(\sum_{a\in\mathcal{A}}\sum_{y\in\{0,1\}}\frac{c_{feat}\beta_{a}+c_{lab}(q_{0\mid a}-p_{0\mid a})}{q_{y\mid a}\varepsilon^{2}}\right)~.$ Moreover, the number of true labels requested by this policy is at most $\widetilde{O}(\sum_{a\in{\mathcal{A}}}\sum_{y\in\{0,1\}}\frac{\beta_{a}}{q_{y\mid a}\varepsilon^{2}})$.
The audit cost and the number of true labels requested by RS-Audit is lower than the baseline. Note that $\frac{\beta_{a}}{q_{y\mid a}}=\frac{\mathbb{P}[f=0,A=a]}{q_{y,a}}\leq\frac{\beta}{q_{y,a}}$. Therefore, the number of true labels requested by the baseline is more than RS-Audit by $\widetilde{O}(\sum_{y\in\{0,1\},a\in{\mathcal{A}}}\frac{\mathbb{P}[f=0,A\neq a]}{q_{y,a}\varepsilon^{2}})$. Similarly, $\frac{q_{0\mid a}-p_{0\mid a}}{q_{y\mid a}}=\frac{\mathbb{P}[f=0,Y=0,A=a]}{q_{y,a}}\leq\frac{\gamma_{00}}{q_{y,a}}$, hence the audit cost of the baseline is more than RS-Audit by $\widetilde{O}(\sum_{y\in\{0,1\},a\in{\mathcal{A}}}\frac{c_{feat}\mathbb{P}[f=0,A\neq a]+c_{lab}\mathbb{P}[f=0,Y=0,A\neq a]}{q_{y,a}\varepsilon^{2}})$. Next, we show a lower bound on the number of true labels requested by any algorithm for auditing (proof in Appendix A.3).
@@
Under the blackbox setting, for any $\varepsilon\in(0,\nicefrac{{1}}{{4}})$ and $\delta\in(0,1)$, there exists a family of instances characterized by $p_{y\mid a}$, $q_{y\mid a}$, and $\beta_{a}$ for $y\in\{0,1\}$ and $a\in{\mathcal{A}}$, even with $\lvert\mathcal{A}\rvert=2$, such that any algorithm with finite expected running time that succeeds in the $(\varepsilon,\delta)$-fairness test must request at least $\widetilde{\Omega}\left(\sum_{a\in\mathcal{A}}\sum_{y\in\{0,1\}}\frac{\beta_{a}}{q_{y\mid a}\varepsilon^{2}}\right)$ true labels in expectation.
Note that the number of true labels requested by RS-Audit (Algorithm 1) is within log factors of the lower bound. This in turn indicates that the characterization of the blackbox model is essentially complete. To enrich our understanding of the fairness audit problem under partial feedback, we consider mixture models under some mild structural assumptions. In the mixture model setting, we make the following assumptions: (i) $\mathbb{P}[f=1\mid Y=y,A=a]\geq\alpha>0$ for some known constant $\alpha$, (ii) ${\mathcal{E}}_{\theta}(X)$ is log-concave in $X$, $\kappa\mathbf{I}\preceq\nabla^{2}W(\theta)\preceq\lambda\mathbf{I}$ with $0<\kappa<\lambda$ and $W$ is $L$-Lipschitz for all $\theta\in\Theta$, $T(X)$ has components polynomial in $X$, $\mathcal{B}(\theta^{*}_{y,a},\frac{1}{\beta})\subset\Theta$ [^1] This assumption is made in LWZ (23) (see their Claim 1) but not explicitly stated. for some $\beta>0$ for all $y\in\{0,1\}$ and $a\in{\mathcal{A}}$. Without loss of generality, let $\kappa\leq 1$ and $L,\beta\geq 1$. All these assumptions except Lipschitzness are from LWZ (23) which are used to recover model parameters from truncated data. The positivity assumption (i) also appears in other partial feedback works like KMC (24) and ensures that enough probability mass lies on the positive classification, allowing the distribution recovery. The assumptions in (ii) ensure computational efficiency and concentration bounds. These assumptions along with Lipschitzness are mild and are satisfied by several popular distributions like Gaussian, Exponential and continuous Bernoulli for bounded parameter sets (see Appendix B.4).
Finally, letting $q_{M,a}=q_{1\mid a}\vee q_{0\mid a}$ and $q_{m,a}=q_{1\mid a}\wedge q_{0\mid a}$, we assume that for all $a\in{\mathcal{A}}$:
$\lVert\theta^{*}_{1,a}-\theta^{*}_{0,a}\rVert\geq\sqrt{\frac{48(L\vee\beta)\beta+3\lambda}{4\kappa\beta^{2}}\log\left(\frac{10q_{M,a}}{q_{m,a}\varepsilon}\right)}$ (1)
This separation assumption between parameters enables an approximate MAP based oracle to learn the mixture weights without requesting labels. Such an assumption has been considered in the context of mixture of spherical Gaussians in Mah (11) and has been shown to be mild (and tight) condition for learning good classifiers (see their Table 1 & Remark 8).
@@
Given $\varepsilon,\delta\in(0,1)$, under the mixture model and the aforementioned assumptions, Exp-Audit (Algorithm 2) succeeds in $(\varepsilon,\delta)$-fairness audit with $\texttt{cost}\leq\widetilde{O}\left(\sum_{a\in{\mathcal{A}}}\left(\frac{c_{feat}\beta_{a}}{q_{m,a}\varepsilon^{2}}+\sum_{y\in\{0,1\}}\frac{c_{lab}(q_{0\mid a}-p_{0\mid a})}{q_{y\mid a}}\right)\right).$
The above theorem shows that under structural assumptions on the data distribution, algorithm with better audit cost exists. Notably, the cost concerning $c_{lab}$ is a problem-dependent constant and independent of $\varepsilon$. Our algorithm consists of two key components. First is to view the historical partial feedback data as a collection of truncated samples, thereby applying algorithms for learning distribution parameters from truncated samples LWZ (23). Second is learning mixture weights with approximate MAP oracle for mixture of exponential family. We believe the first component to be a conceptual contribution since it motivates studying partial feedback, an important setting in fairness in decision-making, through the lens of truncated samples. For the second component, our contribution is a generalization of Mah (11) from spherical Gaussians to exponential family, which we believe could be of independent interest.
In the next sections, we describe RS-Audit and Exp-Audit. Then, we empirically validate our algorithms on the Adult Income BK (96) and Law School WRC (98) datasets.
@@ 4 Fairness Audit under Blackbox Model
Recall that in the audit model under partial feedback, the auditor has to request the true label $Y$ of a negatively classified individual by paying a cost. For the auditor to succeed in fairness audit, it suffices to estimate $\mathbb{P}[f=1\mid Y=y,A=a]$ for all $y\in\{0,1\}$ and $a\in{\mathcal{A}}$. For a given $a\in\mathcal{A}$ and $y\in\{0,1\}$, $\mathbb{P}[f=1\mid Y=y,A=a]=\frac{\mathbb{P}[f=1,Y=y,A=a]}{\mathbb{P}[Y=y,A=a]}=\frac{p_{y,a}}{q_{y,a}}~.$ First, note that $p_{y,a}$ can be estimated from past database using a simple empirical estimate $\frac{1}{n}\sum_{i=1}^{n}\mathds{1}\{f_{i}=1,Y_{i}=y,A_{i}=a\}$ without incurring any cost. As the size of the database $n$ grows, this estimate becomes more accurate. Henceforth, we will focus on estimation of $q_{y,a}$ which requires online data, although we will specify how large $n$ is sufficient for the fairness audit.
@@
Baseline (Algorithm 3). To estimate $q_{y,a}$, a natural policy is to observe the true labels of all negatively classified individuals. To correctly identify the true hypothesis, it is sufficient to estimate $p_{y,a}$ and $q_{y,a}$ within a multiplicative factor of $1\pm\frac{\varepsilon}{12}$ for all $y\in\{0,1\}$ and $a\in{\mathcal{A}}$ (see proof of Theorem 1, Eq. (2) in Appendix A.1). For a fixed $y\in\{0,1\}$ and $a\in{\mathcal{A}}$, $\widetilde{O}(\frac{1}{p_{y,a}\varepsilon^{2}})$ and $\widetilde{O}(\frac{1}{q_{y,a}\varepsilon^{2}})$ samples are sufficient, respectively, via Chernoff bound (Lemma 28). Therefore, the auditor can stop after observing $\widetilde{O}(\frac{1}{q_{y,a}\varepsilon^{2}})$ true labels from online data (including $f=1$ individuals and those the auditor requests). However, this insight does not translate into an algorithm since $p_{y,a}$ and $q_{y,a}$ are unknown.
@@
Hence, we base our algorithm on a random stopping time that depends on an appropriately chosen threshold $\tau$ which is a function of $\varepsilon$, $\delta$ and $\lvert{\mathcal{A}}\rvert$ only. The stopping time for $q_{y,a}$ estimation is defined as $N_{y,a}\coloneqq\min\{t\geq 1:\sum_{i=1}^{t}\mathds{1}\{Y_{i}=y,A_{i}=a\}\geq\tau\}$ and likewise define $N^{\prime}_{y,a}$ for $p_{y,a}$ but with $\mathds{1}\{f_{i}=1,Y_{i}=y,A_{i}=a\}$. In other words, for every $y\in\{0,1\}$ and $a\in{\mathcal{A}}$, the auditor keeps requesting true labels of individuals till a total (including $f=1$ individuals) of $\tau$ individuals with $Y=y$ and $A=a$ are observed. For a $y\in\{0,1\}$ and $a\in{\mathcal{A}}$, we set $\widehat{q}_{y,a}=\frac{\tau}{N_{ya}}$ and $\widehat{p}_{y,a}=\frac{\tau}{N^{\prime}_{y,a}}$ as an estimate of $q_{y,a}$ and $p_{y,a}$ respectively. Using concentration of negative binomials (Lemma 23), we show that these are within $1\pm\frac{\varepsilon}{12}$ factor of $q_{y,a}$ and $p_{y,a}$. The auditor then computes $\widehat{\Delta}=\max_{y\in\{0,1\},a,a^{\prime}\in{\mathcal{A}}}\lvert\frac{\widehat{p}_{y,a}}{\widehat{q}_{y,a}}-\frac{\widehat{p}_{y,a^{\prime}}}{\widehat{q}_{y,a^{\prime}}}\rvert$ and concludes $\mathsf{UNFAIR}$ if $\widehat{\Delta}>\frac{\varepsilon}{2}$ and $\mathsf{FAIR}$ otherwise.

### v2

@@ 3 Our Results
We denote by $q_{y,a}\coloneqq\mathbb{P}[Y=y,A=a]$, $q_{y\mid a}\coloneqq\mathbb{P}[Y=y\mid A=a]$ for all $y\in\{0,1\}$ and $a\in{\mathcal{A}}$. Note that $q_{y,a}=q_{y\mid a}\cdot\mathbb{P}[A=a]$, therefore $q_{y,a}\leq q_{y\mid a}$. Next, we state the key results of this work. All our results hold for arbitrary classifier $f$, and our techniques extend to other fairness metrics like demographic parity and equal opportunity.
@@
Given $\gamma\geq 0$, $\varepsilon,\delta\in(0,1)$, under the blackbox model, the audit policy that requests true labels of every individual with $f=0$ (Algorithm 3 in Appendix A) succeeds in the $(\gamma,\varepsilon,\delta)$-fairness audit with $\texttt{cost}\leq\widetilde{O}\left(\sum_{a\in\mathcal{A}}\sum_{y\in\{0,1\}}\frac{c_{feat}\mathbb{P}[f=0]+c_{lab}\mathbb{P}[f=0,Y=0]}{q_{y,a}\varepsilon^{2}}\right)~.$ Moreover, the number of true labels requested by this policy is at most $\widetilde{O}(\sum_{a\in{\mathcal{A}}}\sum_{y\in\{0,1\}}\frac{\mathbb{P}[f=0]}{q_{y,a}\varepsilon^{2}})$.
Next, we give an algorithm RS-Audit (Algorithm 1) that has improved audit cost using rejection sampling. This algorithm requests an optimal number of true labels, indicating that characterization of the blackbox case is essentially complete. In the design of RS-Audit (and the baseline), we use a threshold-based stopping time construction to decide when to stop requesting true labels. The analysis requires concentration properties of the Negative Binomial distribution, instead of the usual Chernoff bounds (Lemma 28). Below we formally state the guarantees (proof in A.2).
@@
For $\gamma\geq 0$, $\varepsilon,\delta\in(0,1)$, in the blackbox model, Algorithm 1 (RS-Audit) succeeds in $(\gamma,\varepsilon,\delta)$-fairness audit with $\texttt{cost}\leq\widetilde{O}\left(\sum_{a\in\mathcal{A}}\sum_{y\in\{0,1\}}\frac{c_{feat}\mathbb{P}[f=0\mid A=a]+c_{lab}\mathbb{P}[f=0,Y=0\mid A=a]}{q_{y\mid a}\varepsilon^{2}}\right).$ Moreover, the number of true labels requested by this policy is at most $\widetilde{O}(\sum_{a\in{\mathcal{A}}}\sum_{y\in\{0,1\}}\frac{\mathbb{P}[f=0\mid A=a]}{q_{y\mid a}\varepsilon^{2}})$.
RS-Audit runs in polynomial time. The audit cost and the number of true labels requested by it is lower than the baseline since $\frac{\mathbb{P}[f=0,Y=0,A=a]}{q_{y,a}}\leq\frac{\mathbb{P}[f=0,Y=0]}{q_{y,a}}$ and $\frac{\mathbb{P}[f=0,A=a]}{q_{y,a}}\leq\frac{\mathbb{P}[f=0]}{q_{y,a}}$. Hence, the number of true labels requested by the baseline is more than RS-Audit by $\widetilde{\Omega}(\sum_{y\in\{0,1\},a\in{\mathcal{A}}}\frac{\mathbb{P}[f=0,A\neq a]}{q_{y,a}\varepsilon^{2}})$. and the audit cost by $\widetilde{\Omega}(\sum_{y\in\{0,1\},a\in{\mathcal{A}}}\frac{c_{feat}\mathbb{P}[f=0,A\neq a]+c_{lab}\mathbb{P}[f=0,Y=0,A\neq a]}{q_{y,a}\varepsilon^{2}})$. Next, we show a lower bound on the cost incurred by any algorithm for auditing (proof in Appendix A.3).
@@
Under the blackbox setting, for any $\varepsilon\in(0,\nicefrac{{1}}{{8}})$ and $\delta\in(0,1)$, there exists a family of instances characterized by $\gamma=0$, $q_{y\mid a}$, and $\mathbb{P}[f=0\mid A=a]$ for $y\in\{0,1\}$ and $a\in{\mathcal{A}}$, even with $\lvert\mathcal{A}\rvert=2$, such that any algorithm with finite expected running time that succeeds in the $(\varepsilon,\delta)$-fairness test must suffer at least $\widetilde{\Omega}\left(\sum_{a\in\mathcal{A}}\sum_{y\in\{0,1\}}\frac{c_{feat}\mathbb{P}[f=0\mid A=a]+c_{lab}\mathbb{P}[f=0,Y=0\mid A=a]}{q_{y\mid a}\varepsilon^{2}}\right)$ cost in expectation.
Note that the audit cost incurred by RS-Audit (Algorithm 1) is within log factors of the lower bound. This in turn indicates that the characterization of the blackbox model is complete. To enrich our understanding of the fairness audit problem under partial feedback, we consider mixture models under some mild structural assumptions. In the mixture model setting, we make the following assumptions:
Assumption 1 (Structural).
(i) $\mathbb{P}[f=1\mid Y=y,A=a]\geq\alpha>0$ for some known constant $\alpha$, (ii) ${\mathcal{E}}_{\theta}(X)$ is log-concave in $X$, $\kappa\mathbf{I}\preceq\nabla^{2}W(\theta)\preceq\lambda\mathbf{I}$ with $0<\kappa<\lambda$ and $W$ is $L$-Lipschitz for all $\theta\in\Theta$, $T(X)$ has components polynomial in $X$, $\mathcal{B}(\theta^{*}_{y,a},\frac{1}{\beta})\subset\Theta$ for some $\beta>0$ for all $y\in\{0,1\}$ and $a\in{\mathcal{A}}$.
Without loss of generality, let $\kappa\leq 1$ and $L,\beta\geq 1$. All these assumptions except Lipschitzness are from Lee et al. (2023) which are used to recover model parameters from truncated data. The positivity assumption (i) also appears in other partial feedback works like Keswani et al. (2024) and ensures that enough probability mass lies on the positive classification, allowing the distribution recovery. The assumptions in (ii) ensure computational efficiency and concentration bounds. These assumptions along with Lipschitzness are mild and are satisfied by several popular distributions like Gaussian, Exponential and continuous Bernoulli for bounded parameter sets (see Appendix B.4).
Assumption 2 (Separation).
With $q_{M,a}\coloneqq q_{1\mid a}\vee q_{0\mid a}$ and $q_{m,a}\coloneqq q_{1\mid a}\wedge q_{0\mid a}$, let for all $a\in{\mathcal{A}}$:
$\lVert\theta^{*}_{1,a}-\theta^{*}_{0,a}\rVert\geq\sqrt{\frac{48(L\vee\beta)\beta+3\lambda}{8\kappa\beta^{2}}\log\left(\frac{10q_{M,a}}{q_{m,a}\varepsilon}\right)}$ (1)
This separation assumption between parameters enables us to construct an approximate oracle to learn the mixture weights without requesting labels. Such an assumption has been considered in the context of mixture of spherical Gaussians in Mahalanabis (2011) and has been shown to be mild (and tight) condition for low Bayes error rate (see their Table 1 & Remark 8). We provide further discussion in Section 5 Remark 5.
@@
Given $\gamma\geq 0$, $\varepsilon,\delta\in(0,1)$, under the mixture model and assumptions 1 and 2, Exp-Audit (Algorithm 2) succeeds in $(\gamma,\varepsilon,\delta)$-fairness audit with $\texttt{cost}\leq\widetilde{O}\left(\underset{a\in{\mathcal{A}}}{\sum}\left(\frac{c_{feat}\mathbb{P}[f=0\mid A=a]}{q_{m,a}\varepsilon^{2}}+\underset{y\in\{0,1\}}{\sum}\frac{c_{lab}\mathbb{P}[f=0,Y=0\mid A=a]}{q_{y\mid a}}\right)\right).$
Moreover, Exp-Audit runs in polynomial time. The above theorem shows that under structural and separation assumptions on the data distribution, algorithm with better audit cost exists. Notably, the cost term concerning $c_{lab}$ is a problem-dependent constant and independent of $\varepsilon$. Our algorithm consists of two key components. First is to view the historical partial feedback data as a collection of truncated samples, thereby applying algorithms for learning distribution parameters from truncated samples Lee et al. (2023). Second is learning mixture weights with approximate MAP oracle for mixture of exponential family. We believe the first component to be a conceptual contribution since it motivates studying partial feedback, an important setting in fairness in decision-making, through the lens of truncated samples. For the second component, we generalize Mahalanabis (2011) from spherical Gaussians to exponential family, which we believe could be of independent interest.
@@ 4 Fairness Audit under Blackbox Model
Recall that in the audit model under partial feedback, the auditor has to request the true label $Y$ of a negatively classified individual by paying a cost. For the auditor to succeed in fairness audit, it suffices to estimate $\mathbb{P}[f=1\mid Y=y,A=a]$ for all $y\in\{0,1\}$ and $a\in{\mathcal{A}}$. For a given $a\in\mathcal{A}$ and $y\in\{0,1\}$, $\mathbb{P}[f=1\mid Y=y,A=a]=\frac{\mathbb{P}[f=1,Y=y,A=a]}{\mathbb{P}[Y=y,A=a]}.$ First, note that $p_{y,a}\coloneqq\mathbb{P}[f=1,Y=y,A=a]$ can be estimated from past database using a simple empirical estimate $\frac{1}{n}\sum_{i=1}^{n}\mathds{1}\{f_{i}=1,Y_{i}=y,A_{i}=a\}$ without incurring any cost. As the size of the database $n$ grows, this estimate becomes more accurate. Henceforth, we will focus on estimation of $q_{y,a}$ which requires online data, although we will specify how large $n$ is sufficient for the fairness audit.
@@
Baseline (Algorithm 3). To estimate $q_{y,a}$, a natural policy is to observe the true labels of all negatively classified individuals. To correctly identify the true hypothesis, it is sufficient to estimate $p_{y,a}$ and $q_{y,a}$ within a multiplicative factor of $1\pm\frac{\varepsilon}{12}$ for all $y\in\{0,1\}$ and $a\in{\mathcal{A}}$ (see proof of Theorem 1, Eq. (2) in Appendix A.1). For a fixed $y\in\{0,1\}$ and $a\in{\mathcal{A}}$, $\widetilde{O}(\frac{1}{p_{y,a}\varepsilon^{2}})$ and $\widetilde{O}(\frac{1}{q_{y,a}\varepsilon^{2}})$ samples are sufficient, respectively, via Chernoff bound (Lemma 28). Therefore, the auditor can stop after observing $\widetilde{O}(\frac{1}{q_{y,a}\varepsilon^{2}})$ true labels from online data (including $f=1$ individuals, and those the auditor requests). However, this insight does not translate into an algorithm since $p_{y,a}$ and $q_{y,a}$ are unknown.
@@
Hence, we base our algorithm on a random stopping time that depends on an appropriately chosen threshold $\tau$ which is a function of $\varepsilon$, $\delta$ and $\lvert{\mathcal{A}}\rvert$ only. The stopping time for $q_{y,a}$ estimation is defined as $N_{y,a}\coloneqq\min\{t\geq 1:\sum_{i=1}^{t}\mathds{1}\{Y_{i}=y,A_{i}=a\}\geq\tau\}$ and likewise define $N^{\prime}_{y,a}$ for $p_{y,a}$ but with $\mathds{1}\{f_{i}=1,Y_{i}=y,A_{i}=a\}$. In other words, for every $y\in\{0,1\}$ and $a\in{\mathcal{A}}$, the auditor keeps requesting true labels of individuals till a total (including $f=1$ individuals) of $\tau$ individuals with $Y=y$ and $A=a$ are observed. Thereafter, $\widehat{q}_{y,a}=\frac{\tau}{N_{y,a}}$ and $\widehat{p}_{y,a}=\frac{\tau}{N^{\prime}_{y,a}}$ act as an estimate of $q_{y,a}$ and $p_{y,a}$ respectively. Using concentration of Negative Binomials (Lemma 23), we show that these estimates are within $1\pm\frac{\varepsilon}{12}$ factor of $q_{y,a}$ and $p_{y,a}$. The auditor then computes $\widehat{\Delta}=\max_{y\in\{0,1\},a,a^{\prime}\in{\mathcal{A}}}\lvert\frac{\widehat{p}_{y,a}}{\widehat{q}_{y,a}}-\frac{\widehat{p}_{y,a^{\prime}}}{\widehat{q}_{y,a^{\prime}}}\rvert$ and concludes $\mathsf{UNFAIR}$ if $\widehat{\Delta}>\gamma+\frac{\varepsilon}{2}$ and $\mathsf{FAIR}$ otherwise.